OpenStudy (b77w):

It has been found that times taken by people to complete a particular tax form follow a normal distribution with mean 100 minutes and standard deviation 30 minutes. A random sample of nine people who have completed this tax form was taken. What is the probability that the sample mean time taken is more than two hours?

jimthompson5910 (jim_thompson5910):

how far did you get with this?

OpenStudy (b77w):

I started calculating something expected values or something but i cant seem to find a formula for probability and samples

jimthompson5910 (jim_thompson5910):

first thing to do is rewrite 2 hrs as 120 minutes

jimthompson5910 (jim_thompson5910):

we're going to be working with the sample mean distribution (basically the distribution of xbar) so that tells us the mean = mu standard deviation = sigma/sqrt(n) hopefully this looks familiar?

OpenStudy (b77w):

Yeah i have that

jimthompson5910 (jim_thompson5910):

ok what is the standard deviation of this particular sample mean distribution?

jimthompson5910 (jim_thompson5910):

\[\Large SE = \frac{\sigma}{\sqrt{n}} = \frac{30}{\sqrt{9}} = ??\]

OpenStudy (b77w):

10